OpenStudy (anonymous):

write the equation in the form axsquared+bx+c=0: (2/x)squared-4=7

OpenStudy (lgbasallote):

\[(\frac{2}{x})^2 - 4 = 7\] \[\frac{4}{x^2} - 4 - 7 = 0\] \[\frac{4}{x^2} - 11 = 0\] \[4 - 11x^2 = 0\] got that?

OpenStudy (lgbasallote):

actually you have to rearrange it \[-11x^2 + 4 = 0\] \[11x^2 - 4 = 0\] so any unclear steps?

OpenStudy (anonymous):

but...it has no b..., there is just the a=11 & the c=-4

OpenStudy (lgbasallote):

yup b = 0

OpenStudy (anonymous):

ah...... ok its clear 2 me now

OpenStudy (anonymous):

tnx alot

OpenStudy (lgbasallote):

great! that's nice to hear :)